Conundrum, Inc. manufactures puzzles. Due to a fire in the administrative offices, the accounting records for September of the current year were partially destroyed. You have been able to piece together the following information from the ledger.


Raw-Material Inventory Accounts Payable

                                    40,000

  12,000                   Bal. 8/31

   Bal. 9/30                    45,000

                  

                   

  


Work-in-Process Inventory Finished-Goods Inventory

   Bal. 8/31                      8,000

   

   Bal. 8/31                    35,000


Manufacturing Overhead Cost of Goods Sold

                                     60,000  

  

  


Wages Payable Sales Revenue

   1,700                     Bal 9/30


Accounts Receivable

   Bal. 8/31                     8,000


Upon examining various source documents and interviewing several employees, you were able to gather the following additional information.


a. Collections of accounts receivable during September amounted to $205,000.
b. Sales revenue in September was 120 percent of cost of goods sold. All sales are on account.
c.

Overhead is applied using an annual predetermined overhead rate using direct-labor hours (based on practical capacity).

d. The budgeted overhead for the current year is $720,000.
e.

Budgeted direct-labor cost for the current year is $960,000. The direct-labor rate is $20 per hour.

f.

The accounts payable balance on September 30 was $1,000. Only purchases of raw material are credited to accounts payable. A payment of $81,000 was made on September 15.

g. September’s cost of goods sold amounted to $182,000.
h. The September 30 balance in finished-goods inventory was $3,250.
i.

Payments of $79,500 were made to direct-labor employees during September. The August 31 balance in the Wages Payable account was $1,000.

j. The actual manufacturing overhead for September was $60,000.
k.

An analysis of the puzzles still in process on September 30 revealed that so far these items have required 500 hours of direct labor and $20,600 of direct material.

Required: a. Calculate the following amounts. 1. Sales revenue for September. 2 September 30 balance in accounts receivable 3. Cost of raw material purchased during September 4. September 30 balance in work-in-process inventory 5. Direct labor added to work in process during September 6. Applied overhead for September. 7. Cost of goods completed during September 8. Raw material used during September. 9. August 31 balance in raw-material inventory. 10. Overapplied or underapplied overhead for September.

b. Complete the T-accounts with the help of the additional information given in the problem. Raw-Material Inventory Accounts Payable 12,000 Bal. 8/31 Bal. 8/31 40,000 Bal. 9/30 Bal. 9/30 Work-in-Process Inventory Finished-Goods Inventory Bal. 8/31 Direct material Direct labor Overhead 8,000 Bal. 8/31 35,000 Bal. 9/30 Bal. 9/30 Manufacturing Overhead Cost of Goods Sold 60,000 Wages Payable Sales Revenue Bal. 8/31 Bal. 9/30 Accounts Receivable Bal. 8/31 8,000 Bal. 9/30
Answer #1

a.

1 Sales revenue for September = 120% x $182000 = $218400
2 September 30 balance in accounts receivable = $8000 + 218400 - 205000 = $21400
3 Cost of raw material purchased during September = $1000 + 81000 - 12000 = $70000
4 September 30 balance in work-in-process inventory = $20600 + (500 x $20) + (500 x $15) = $20600 + 10000 + 7500 = $38100
5 Direct labor added to work in process during September = $79500 + 1700 - 1000 = $80200
6 Applied overhead for September = 4010 x $15 = $60150
7 Cost of goods completed during September = $182000 + 3250 - 35000 = $150250
8 Raw material used during September = $40000
9 August 31 balance in raw material inventory = $40000 + 45000 - 70000 = $15000
10 Overapplied or underapplied overhead for September = $60150 - 60000 = $150 overapplied

Working:

Predetermined overhead rate = Budgeted overhead/Budgeted direct labor hours = $720000/48000 = $15 per direct labor hour

Budgeted direct labor hours = Budgeted direct labor cost/Direct labor rate = $960000/$20 = 48000

Actual direct labor hours during September = $80200/$20 = 4010

b.

Raw Material Inventory Accounts Payable
Bal. 8/31 15000 Bal. 8/31 12000
3 70000 40000 8 f. 81000 70000 3
Bal. 9/30 45000 Bal. 9/30 1000 f.
Work-in-Process Inventory Finished Goods Inventory
Bal. 8/31 8000 Bal. 8/31 35000
Direct
material
40000 7 150250 182000 g.
Direct
labor
80200
Overhead 60150
150250 7
Bal. 9/30 38100 Bal. 9/30 3250
Manufacturing Overhead Cost of Goods Sold
j. 60000 60150 g. 182000
Wages Payable Sales Revenue
Bal. 8/31 1000 218400 1
i. 79500 80200
Bal. 9/30 1700
Accounts Receivable
Bal. 8/31 8000
1 218400 205000 a.
Bal. 9/30 21400